Solva for Z: |6z-3| + 5 I go z as equal to 0 and 1
Answer by Alan3354(69443)

Solva for Z: |6z-3| + 5 I go z as equal to 0 and 1
-----------------
|6z-3| + 5 <= 8
|6z+3| <= 3
----------
Case 1:
6z + 3 <= 3
6z <= 0
z <= 0
--------
Case 2:
-6z-3 <= 3
-6z <= 6
z >= -1
---------
The solution can't be "equal", it's not an equation.
z can be any value between -1 and zero, inclusive
{z| -1 =< z =< 0}